OpenStudy (anonymous):

Seven slips of paper marked with the numbers 1, 2, 3, 4, 5, 6, and 7 are placed in a box and mixed well. Two are drawn. What are the odds that the sum of the numbers on the two slips is 8?

OpenStudy (turingtest):

how many ways can you combine these numbers to add to 8 ?

OpenStudy (turingtest):

how many ways can you pull out 7 slips of paper 2 at a time?

OpenStudy (anonymous):

A 1 to 7 B 1 to 9 C 1 to 6 D 2 to 21

OpenStudy (anonymous):

I'm confused, @TuringTest what are explaining?

OpenStudy (turingtest):

I don't understand your multiple choices o-0 I was trying to find the total number of ways to make 8 from two of those numbers divided by the total number of ways to draw the two papers

OpenStudy (anonymous):

It's not asking for the number of ways. It's asking the odds

OpenStudy (turingtest):

at least I'm not getting any of those as an answer is what I mean, I must have made a mistake

OpenStudy (anonymous):

im confused completely...

OpenStudy (turingtest):

oh my mistake, only two at a time... there are 1+7 2+6 3+5 ---- 3 ways to make 8 and there are 7 ways to draw the first number, and 6 ways to draw the second ways to make 8 ------------- = 3/(7*6)=1/14 total ways to draw hm.. not an answer choice I'm sorry I guess I'm bad at probablility

OpenStudy (turingtest):

I'm bad at spelling it too

OpenStudy (anonymous):

hahah its okay! Thanks though:)

OpenStudy (turingtest):

good luck next time!

OpenStudy (turingtest):

oh wait, if the order of drawing them matter then there are twice as many ways to make 8 that is, if 1+7 is distinct from 7+1 that would double the number of ways to make 8, which \(is\) an answer choice